calculate the solubility of Mg(OH) in moles per liter if the solubility product of Mg(OH)2 is 5 x 10^12
1
Expert's answer
2020-12-16T05:52:22-0500

Mg(OH)2 ------------- Mg2+ +2(OH)-

Solubility product=[Mg2+][OH-]2

Let the conc of Mg2+= x

So (OH-) =2x

So we have;

5×1012=(x)(2x)2

5×1012=4x3

x3=1.25×1012

x=1.08×104

Therefore the Solubility of Mg(OH)2=1.08×104